OpenStudy (anonymous):

To measure the height of one of the faces on Mt. Rushmore, two sightings 810 feet from the base of the mountain are taken. If the angle of elevation to the bottom of the face is 32° and the angle of elevation to the top is 35°, what is the height of this face? (answer to the nearest foot)

OpenStudy (amistre64):

you subtract tangents

OpenStudy (amistre64):

tan(32) = h/810 tan(35) = H/810 H-h = face height :)

OpenStudy (amistre64):

810 [tan(35) - tan(32)] = face height

OpenStudy (anonymous):

okay. so the height in feet is 61.02 feet?

OpenStudy (amistre64):

thats what the calulator i have tells me :)

OpenStudy (anonymous):

haha okay thanks!